Describe in words the graph of each of these curves below. Include in your description the shape, along with other possible relevant information such as length, width, and center points.
a. Y = 3X2
b. (X-1)2 + (Y-8)2 = 16
c. (X+2)2 + (Y-4)2 = 36
d. Y = X2 - X

Answer by stanbon(48558)

Describe in words the graph of each of these curves below. Include in your description the shape, along with other possible relevant information such as length, width, and center points.
a. Y = 3X^2
parabola opening upward with vertex at (0,0)
---------------
b. (X-1)^2 + (Y-8)^2 = 16
circle with center at (1,8) and radius = 4
---------------------
c. (X+2)^2 + (Y-4)^2 = 36
circle with center at (-2,4) and radius = 6
---------------------
d. Y = X^2 - X
parabola opening upward with vertex at (1/2, -1/4)
===============
Cheers,
Stan H.